Fermi National Laboratory
 

Accelerator Update for February 3, 2006 - February 6, 2006

 

Friday February 3

The day shift began with Operations monitoring the store 4621 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with Switchyard (SY), NuMI, and MiniBooNE taking beam.

Operations terminated store 4621 at 8:15 AM.

Operations prepared D-Zero (D0) for a 1.5-hour collision hall access at 8:35 AM.

Operations put the Tevatron (TeV) through a dry squeeze at 10:01 AM and then put the TeV into shot setup.

Operations established store 4623 at 1:02 PM with an initial luminosity of 99.47E30.

The Recycler began stashing antiprotons at 1:05 PM. The stash was completed around 2 PM.

At 1:45 PM, Operations turned off the Booster RF so experts could repair a RF station (BRF4). The repair took an hour.

Pbar experts accessed their tunnel to repair a quadrupole shunt (QS732) at 2:28 PM.

Pbar resumed stacking at 5:19 PM.

At 9:09 PM, the Tevatron Electron Lens (TEL) tripped off for the first of many times during the evening shift.

The Recycler began stashing at 9:25 PM. The stash ended and Pbar resumed stacking at 10:14 PM.

At 11:57 PM, a quadrupole (I:Q703) power supply in the P1 transfer line tripped off and tripped its yard breaker. The Duty Electrician was informed.

Saturday February 4

The midnight shift began with Operations monitoring store 4623 and the stash, and with beam off due to a problem with Q703. Experts began working at resolving all the TEL trips.

At 12:48 AM, Operations reported that a power supply expert was called in to investigate Q703. NUMI was the only experiment taking beam.

Pbar resumed stacking at 3:01 AM.

TEL experts reported at 3:30 AM that they temporarily moved the remote ON/OFF switch outside of the chassis. The TEL had remained on for an hour without tripping.

At 4:35 AM, the Booster East Anode Power Supply (EAPS) tripped off. Operators investigated. At 5 AM they reported that BRF5 also tripped off and would not reset. Operators bypassed the station.

The Recycler stashed at 5:16 AM. The stash ended and MCenter resumed taking beam at 5:59 AM. Pbar resumed stacking at 6:04 AM.

The Recycler began stashing at 9:18 AM. The stash ended and MCenter and MTest resumed taking beam at 10:54 AM. Pbar resumed stacking.

At 11:06 AM, Operations shut off the Booster beam so an RF expert could make repairs on BRF5. The expert completed his repairs and beam resumed at 12:03 PM.

Operations terminated store 4623 at 3:59 PM.

Operations put the TeV into shot setup at 4:13 PM.

Operations established store 4624 at 6:08 PM with an initial luminosity of 104.78E30.

The Recycler began preparing to stash antiprotons at 6:31 PM. The stash ended at 6:52 PM. Pbar, MTest, and MCenter resumed taking beam.

At 7:07 PM, the TeV suffered a sector F4 quench. Operators investigated and then called in controls experts. The experts discovered a CAMAC crate with a flakey power supply.

Sunday February 5

The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash, with Pbar stacking, and with MTest, MCenter, NuMI, and MiniBooNE taking beam.

The Recycler began preparing to stash at 1:27 AM.

Operations put the TeV through a dry squeeze at 1:32 AM.

The MTest experiment shut down at 1:34 AM.

The stash ended and Pbar resumed stacking at 2:17 AM. MCenter resumed taking beam.

Operations put the TeV through a wet squeeze at 2:38 AM.

At 3:12 AM, the NuMI Horn tripped off. Operators contacted an expert.

Operations put the TeV into shot setup at 4:11 AM.

Operations established store 4627 at 5:51 AM with an initial luminosity of 85.7E30. Pbar resumed stacking.

The Recycler began preparing to stash at 7:25 AM. The stash ended and MCenter resumed taking beam at 8:19 AM.

MTest resumed taking beam at 8:23 AM.

Pbar resumed stacking at 8:24 AM.

At 10:57 AM, Operations disabled beam to MTest and MCenter so that Pbar experts could begin a Pbar/NuMI study period. The study ended at 3:33 PM.

Pbar, MTest, and MCenter resumed taking beam by 4:11 PM.

The Recycler began preparing to stash at 7:39 PM. The stash ended around 8 PM. Pbar, MTest, and MCenter resumed taking beam.

Monday February 6

The midnight shift began with Operations monitoring store 4627 and the stash, with Pbar stacking, and with MTest, MCenter, NuMI, and MiniBooNE taking beam.

The Recycler began preparing to stash at 12:56 AM. The stash ended around 1:42 AM. Pbar and MCenter resumed taking beam. MTest shut down for the night.

At 3:21 AM, a Booster kicker (MKS08) tripped off. Operators reset the device, but it tripped off again. Operators called in experts to investigate.

At 4:55 AM, an Accumulator quadrupole power supply (A:LQ) tripped off on a ground fault indication. The stack was lost due to this trip. Operations called in an expert.

Booster experts had the kicker fixed by 5:20 AM. Beam to MiniBooNE, NuMI, and SY resumed.

Operations terminated store 4727 at 5:52 AM.

Operations put the TeV into shot setup at 6:05 AM.

Operations prepared the Pbar tunnel for an access at 6:53 AM.

The Plans for Monday

The plans are to establish a store early this morning and to repair the Pbar quadrupole.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/06/2006    email Fermilab